OpenStudy (anonymous):

If z=1/2 what is the value of -(z^2z^-4)^-3

Directrix (directrix):

- ( z^2 z^-4 )^-3 = - [ (1/2)² z ^ (-4) ] ^ (-3)

OpenStudy (mertsj):

-z^6=-(1/2)^6=-1/64

Directrix (directrix):

@bih 122 --> Would you like me to continue with the steps?

OpenStudy (anonymous):

yes please

Directrix (directrix):

It is easier to write it given that this site is slow tonight.

Directrix (directrix):

I used the power of powers multiplication rule to distribute the exponent -3. Questions?

OpenStudy (anonymous):

how did the other guy get -1/64

Directrix (directrix):

We're not finished. Continuing, -z ^ 12 / z^6 = - z ^ 6 = - (1/2) ^ 6 = and ...

Directrix (directrix):

- (1/2) ^ 6 = - [ 1 / 2^6] = - 1 / 64 ---> Answer

OpenStudy (anonymous):

o ok thanks

Directrix (directrix):

Any time.